OpenStudy (anonymous):

Easy but cute question ,show that:- \(\sum _{k=0}^n 10^k=\frac{10^{n+1}-1}{9}\)

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

Induction: For \(n=0\), you have \[\begin{align*}10^0=1&=\frac{10^1-1}{9}\\ &\stackrel{\checkmark}{=}1 \end{align*}\] Assume \[\sum_{k=0}^n10^k=\frac{10^{n+1}-1}{9}\] Then \[\begin{align*}\sum_{k=0}^{n+1}10^k&=\sum_{k=0}^{n}10^k+10^{n+1}\\\\ &=\frac{10^{n+1}-1}{9}+\frac{9\left(10^{n+1}\right)}{9}\\\\ &=\frac{10\left(10^{n+1}\right)-1}{9}\\\\ &=\frac{10^{n+2}-1}{9}\end{align*}\] and we're done :)

OpenStudy (anonymous):

nice